# An exact approach for the multi-depot electric vehicle scheduling problem

### v1

@@ 6.3 Results for a low number of depots
Table 3: Results for all formulations for a low number of depots
setting #opt $t$(s) gap(%) #nB&B $z^{*}$ $LB_{R}$cuts $t_{cut}$(s)
3i+VI 176 3,824.48 0.620 431,497.1 1,254,985.23 1,229,348.96
2i-IP+VI I One 191 3,463.46 0.530 567,248.1 1,252,516.59 1,226,759.39 632.23 6.14
I All 202 2,860.34 0.451 538,242.6 1,254,628.51 1,228,940.49 382.85 1.27
IF One 193 3,498.09 0.582 197,146.9 1,255,224.46 1,229,290.50 157,296.92 1,236.79
IF All 189 3,499.67 0.589 179,160.5 1,255,218.01 1,229,209.67 454,343.61 1,354.13
2i-CC+VI I One 194 3,406.85 0.518 490,587.1 1,255,123.02 1,229,223.12 379.12 6.70
I All 198 3,206.57 0.537 438,556.6 1,255,092.29 1,229,179.49 262.35 4.27
IF One 137 5,594.86 10.158 5,604.3 1,240,238.23 1,229,220.04 35,474.35 293.37
IF All 142 5,376.71 9.888 6,043.3 1,237,369.33 1,229,213.33 31,963.85 221.43
@@
  • the 3-index formulation (1) (named 3i) with (VI) valid inequalities  (4a), (5a), (5b) and (5c),
  • the 2-index formulation (2), which includes the infeasible path constraints (2f) (named 2i-IP), with separating only integer (I) or both integer and fractional (IF) solutions, with including all (All) or only the first (One) infeasible path constraint, with (VI) valid inequalities (4b), (6a), (6b) and (6c), and
  • the 2-index formulation (2) without (2f) but with the connectivity constraints (3) (named 2i-CC), with separating only integer (I) or both integer and fractional (IF) solutions, with including all (All) or only the first (One) connectivity constraint, and with (VI) valid inequalities (4b), (6a), (6b) and (6c).
For each of these settings, we give the number (#opt) of instances that were solved to optimality (out of 270), the average CPU times ($t$) in seconds, the average percentage gap at the time limit (3 hours) of all instances that were not solved to optimality, the average number of B&B nodes (#nB&B), the average best found objective function value $z^{*}$ at termination, the average lower bound at the root node $LB_{R}$, the average number of cuts (i.e., added infeasible path constraints (2f) or added connectivity constraints (3)) and the average CPU time $t_{cut}$ in seconds for the separation. For instances where no feasible solution was found within the time limit, we assume a gap of 100% and exclude $z^{*}$ from the calculations.
The results in Table 3 clearly show that for 2i-CC+VI all versions with separating fractional and integer solutions IF perform much worse and solve much fewer instances to optimality within the time limit, demonstrating the superiority of I for 2i-CC+VI.
For 2i-IP+VI the picture is not that clear, as in the setting One, IF performs slightly better than I with two more solved instances (193 vs. 191), while in the setting All, IF solves less instances than I (189 vs. 202).
Furthermore, the 2-index formulation with infeasible path constraints (2i-IP+VI) and both I and IF separations outperform the 3-index model in all settings (at least 189 instances are solved to optimality in all 2i-IP+VI settings, while 176 instances are solved to optimality in 3i+VI). Overall, the best setting of this computational evaluation is 2i-IP+VI+I+All with 202 instances solved to optimality.
@@ 6.4 Results for a high number of depots
Table 4: Results for 2-index formulations for a high number of depots
#opt $t$(s) gap(%) #nB&B $z^{*}$ $LB_{R}$cuts $t_{cut}$(s)
2i-IP+VI+I One 44 7,730.64 51.28 591,526.70 1,532,739.78 1,736,747.63 6,899.24 102.15
2i-IP+VI+I All 80 6,123.33 0.60 670,666.60 1,776,004.59 1,740,636.52 3,483.97 21.45
2i-CC+VI+I One 52 6,883.98 23.58 420,146.80 1,684,375.49 1,737,303.19 2,653.12 108.58
2i-CC+VI+I All 55 7,135.10 8.61 398,538.90 1,749,854.24 1,738,969.85 1,593.65 36.80
@@
The settings 2i-IP+VI+I and 2i-CC+VI+I in combination with adding All cuts solve 80 and 55 instances out of 180 to optimality, respectively, and clearly outperform both settings with One. When comparing computation times, 2i-IP+VI+I+All requires 6,123.33 seconds on average and solves 80 instances to optimality, while 2i-CC+VI+I+All requires 7,135.10 seconds on average and solves 55 instances to optimality, indicating that 2i-IP+VI+I+All has a clear edge over 2i-CC. Overall, not only the number of optimally solved instances, but also the optimality gap is much better for both settings with All cuts added. In the end, the setting 2i-IP+VI+I+All can be determined as clear winner in these runs.
@@
Table 5: Detailed results of 2-index and 3-index formulation for a high number of depots
2-index (2i-IP+VI+I+All) 3-index (3i+VI)
$|V^{I}|$ $|K|$opt gap(%) $t$(s) #opt gap(%) $t$(s)
30 6 14 0.32 1,915.68 11 0.58 3,605.33
30 8 15 0.00 21.43 14 0.88 1,246.07
40 6 12 0.54 3,241.20 7 0.63 6,151.67
40 8 15 0.00 895.09 6 0.57 6,307.18
50 6 5 0.45 7,779.66 1 0.80 10,051.77
50 8 10 0.46 4,441.73 2 0.84 9,292.28
60 6 2 0.39 10,210.79 0 1.38 10,804.36
60 8 3 0.52 9,860.92 0 1.42 10,826.87
70 6 1 0.49 10,806.17 0 1.88 10,807.44
70 8 1 0.69 10,362.42 0 2.02 10,808.52
80 6 0 0.87 10,820.71 0 2.35 10,821.09
80 8 2 0.85 9,709.50 0 2.16 10,815.70

### v2

@@ 6.3 Results for a low number of depots
Table 3: Results for all formulations for a low number of depots.
setting #opt tt(s) gap(%) #nB&B $z^{*}$ $LB_{R}$ #cuts $t_{cut}$(s)
3i+VI 175 4,287.84 0.22 400,084.60 1,265,436.67 1,232,878.84
2i-IP+VI I One 196 3,353.96 0.14 529,983.00 1,265,443.29 1,232,935.81 498.80 5.08
I All 206 2,966.02 0.12 443,823.40 1,265,347.58 1,232,880.87 299.49 1.77
IF One 199 3,183.39 0.15 163,507.80 1,265,500.31 1,232,976.81 122,771.76 1,001.90
IF All 196 3,392.24 0.53 162,428.40 1,265,598.42 1,232,796.79 364,862.30 1,262.44
2i-CC+VI I One 202 3,089.69 0.14 412,167.30 1,265,398.67 1,232,943.92 323.79 6.16
I All 206 2,995.78 0.14 385,027.70 1,265,393.44 1,232,887.54 226.26 4.07
IF One 141 5,423.97 4.84 3,974.10 1,248,703.81 1,233,114.30 30,155.29 302.43
IF All 141 5,335.53 4.88 4,677.80 1,245,562.88 1,233,046.19 25,999.09 189.79
@@
  • the 3-index formulation (1) (named 3i) with (VI) valid inequalities (4a), (5a), (5b), and (5c),
  • the 2-index formulation (2), which includes the infeasible path constraints (2f) (named 2i-IP), with separating only integer (I) or both integer and fractional (IF) solutions, with including all (All) or only the first (One) infeasible path constraint, with (VI) valid inequalities (4b), (6a), (6b), and (6c), and
  • the 2-index formulation (2) without (2f) but with the connectivity constraints (3) (named 2i-CC), with separating only integer (I) or both integer and fractional (IF) solutions, with including all (All) or only the first (One) connectivity constraint, and with (VI) valid inequalities (4b), (6a), (6b), and (6c).
For each of these settings, we give the number (#opt) of instances that were solved to optimality (out of 270), the average CPU times ($t$) in seconds, the average percentage gap at the time limit (3 hours) of all instances, the average number of B&B nodes (#nB&B), the average best found objective function value ($z^{*}$) at termination, the average lower bound at the root node ($LB_{R}$), the average number of cuts (#cuts, i.e., added infeasible path constraints (2f) or added connectivity constraints (3)) and the average CPU time ($t_{cut}$) in seconds for the separation. For instances where no feasible solution was found within the time limit, we assume a gap of 100% and exclude $z^{*}$ from the calculations.
The results in Table 3 clearly show that for 2i-CC+VI both versions separating fractional and integer solutions IF perform much worse and solve much fewer instances to optimality within the time limit, demonstrating the superiority of I for 2i-CC+VI.
For 2i-IP+VI the picture is not that clear, as in the setting One, IF performs slightly better than I with three more solved instances (199 vs. 196), while in the setting All, IF solves less instances than I (196 vs. 206).
Furthermore, the 2-index formulation with infeasible path constraints (2i-IP+VI) and both I and IF separations, as well as the 2-index formulation with connectivity cuts (2i-CC+VI) and I separations, outperform the 3-index model in all settings (at least 196 instances are solved to optimality in all 2i-I+VI settings, while 175 instances are solved to optimality in 3i+VI). Overall, the best settings of this computational evaluation are 2i-IP+VI+I+All and 2i-CC+VI+I+All with 206 instances solved to optimality each.
@@ 6.4 Results for a high number of depots
Table 4: Results for 2-index formulations and a high number of depots.
#opt tt(s) gap(%) #nB&B $z^{*}$ $LB_{R}$ #cuts $t_{cut}$(s)
2i-IP+VI+I One 46 8,420.18 38.02 536,048.60 1,579,065.89 1,785,904.08 6,224.49 92.52
2i-IP+VI+I All 60 7,478.90 14.44 408,568.40 1,779,073.37 1,785,904.24 2,903.48 7.20
2i-CC+VI+I One 58 7,601.72 14.95 339,778.70 1,768,086.18 1,785,903.89 2,352.59 72.59
2i-CC+VI+I All 65 7,308.79 4.92 300,350.80 1,832,566.14 1,785,905.88 1,474.92 24.26
@@
The settings 2i-IP+VI+I and 2i-CC+VI+I in combination with adding All cuts solve 60 and 65 instances out of 180 to optimality, respectively, and clearly outperform both settings with One. When comparing computation times, 2i-IP+VI+I+All requires 7,478.90 seconds on average and solves 60 instances to optimality, while 2i-CC+VI+I+All requires 7,308.79 seconds on average and solves 65 instances to optimality, indicating that 2i-CC+VI+I+All has a clear edge over 2i-IP for instances with a high number of depots. Overall, not only the number of optimally solved instances, but also the optimality gap is much better for both settings with All cuts added. In the end, the setting 2i-CC+VI+I+All can be determined as the clear winner in these runs.
@@
Table 5: Detailed results of the 2-index and 3-index formulations for a high number of depots.
2-index (2i-IP+VI+I+All) 2-index (2i-CC+VI+I+All) 3-index (3i+VI)
$|V^{I}|$ $|K|$ #opt tt(s) gap(%) #opt tt(s) gap(%) #opt tt(s) gap(%)
30 6 15 311.64 0.00 15 142.63 0.00 9 4,617.07 0.23
30 8 15 312.54 0.00 15 113.45 0.00 10 4,143.91 0.35
40 6 10 4,088.86 0.12 11 3,540.53 0.07 6 6,830.38 0.43
40 8 11 3,912.17 0.08 11 4,199.32 0.08 4 8,433.61 0.88
50 6 6 6,654.56 0.44 6 6,599.81 0.45 4 9,272.83 1.04
50 8 3 9,573.76 0.54 6 8,441.46 0.43 1 10,707.20 1.81
60 6 0 10,803.65 7.34 0 10,803.51 0.84 0 10,803.86 1.36
60 8 0 10,803.75 1.07 1 10,622.99 0.93 0 10,803.65 1.95